# Random Sperner lemma and random Brouwer fixed point theorem

### v3

@@ 6 The equivalence of random Brouwer fixed point theorem and random Borsuk theorem
It is easy to see that both $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ and $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ are $\sigma$-stable. If $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is a $\mathcal{T}_{c}$-$r$-image of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, then there are a $\sigma$-stable $\mathcal{T}_{c}$-continuous mapping $h:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ and a $\sigma$-stable $\mathcal{T}_{c}$-continuous mapping $g:L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $h\circ g$ is the identity on $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$. Define $f:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ by $f(x)=g(-h(x))$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, then $f$ is a $\sigma$-stable and $\mathcal{T}_{c}$-continuous mapping, and hence there exists $z\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ by Theorem 2.1 such that $f(z)=z$, that is to say, $z=g(-h(z))$, so $h(z)=(h\circ g)(-h(z))=-h(z)$. Thus $h(z)=0\in L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$, a contradiction. ∎
Corollary 6.5 below is essentially Lemma 3.5 of [11].
@@
Theorem 6.4 is equivalent to random Brouwer fixed point theorem for the random closed unite ball $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (and hence also equivalent to Theorem 2.1 by Proposition 6.1).
@@
Suppose that there exists a $\sigma$-stable and $\mathcal{T}_{c}$-continuous mapping $f:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $f$ has no fixed point. Define $e:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F})$ by $e(x)=\|f(x)-x\|$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, and denote $\eta=\bigwedge\{e(x):x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\}$, then by Lemma 3.4 of [26] there exists $x_{0}\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $\eta=\|f(x_{0})-x_{0}\|$ since $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\}$ is random sequentially compact and $\sigma$-stable and since $e$ is $\sigma$-stable and random sequentially continuous. Then $\eta>0$, since otherwise, $\eta=0$ yields $f(x_{0})=x_{0}$, which contradicts the assumption that $f$ has no fixed point.
Now, let $A\in\mathcal{F}$ be an arbitrary chosen representative of $(\eta>0)$, then $P(A)>0$. Since $\theta\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ and $f:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ is $\sigma$-stable, then $f$ has the local property, namely $\tilde{I}_{C}f(x)=\tilde{I}_{C}f(\tilde{I}_{C}x)$ for any $C\in\mathcal{F}$ and $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$: in fact, since
@@
then $\tilde{I}_{C}f(\tilde{I}_{C}x)=\tilde{I}_{C}f(x)$. In particular, $\tilde{I}_{A}f(\tilde{I}_{A}x)=\tilde{I}_{A}f(x)$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. Further, let
@@
and define $f_{A}:G_{A}\rightarrow G_{A}$ by $f_{A}(\tilde{I}_{A}x)=\tilde{I}_{A}f(x)$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, then $f_{A}$ is well defined. It is clear that $G_{A}$ is $\sigma$-stable, $\mathcal{T}_{c}$-closed, a.s. bounded and $L^{0}(\mathcal{F})$-convex subset of $L^{0}(\mathcal{F},R^{d})$ and $f_{A}$ is also $\sigma$-stable and $\mathcal{T}_{c}$-continuous.
Let $(A,A\cap\mathcal{F},P_{A})$ be the probability space with $P_{A}:A\cap\mathcal{F}\rightarrow[0,1]$ defined by $P_{A}(A\cap C)=\frac{P(A\cap C)}{P(A)}$ for any $C\in\mathcal{F}$. Further, identifying $G_{A}$ with $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$, where $\mathcal{F}_{A}=A\cap\mathcal{F}$, then $f_{A}$ can be also identified with a mapping from $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$ to $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$. Again observing $\|f(x)-x\|>0$ on $A$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, namely $\|f_{A}(x)-x\|>0$ on $A$ for each $x\in G_{A}$, we can, without loss of generality, assume that $A=\Omega$, then $\|f(x)-x\|>0$ on $\Omega$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ (since otherwise, we can consider $f_{A}$ and $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$ to replace $f$ and $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$).
Finally, let $L_{x}=\{\lambda\in L^{0}(\mathcal{F}):\|x+\lambda(f(x)-x)\|=1\}$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ and $\lambda(x)=Min~{}L_{x}:=$ the smallest element of $L_{x}$. In fact, $L_{x}$ has only the two elements, which satisfy the equation
@@
where $\langle x,y\rangle$ stands for the $L^{0}$-inner product of $x$ and $y$ in $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Similar to the Vieta theorem, one can have
@@
This define a mapping $\lambda:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F})$ and it is not very difficult to verify that $\lambda$ is still $\sigma$-stable and $\mathcal{T}_{c}$-continuous. By noticing that $\langle x,x-f(x)\rangle=\|x\|^{2}-\langle x,f(x)\rangle$ and the random Schwartz inequality: $|\langle x,f(x)\rangle|\leq\|x\|\|f(x)\|$, then
@@
for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. In particular, when $x\in L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$, namely when $\|x\|=1$, one has $\lambda(x)=0$. Again, define $h:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ by $h(x)=x+\lambda(x)(f(x)-x)$, then $h$ is a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ onto $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$, which contradicts Theorem 6.4. ∎
@@
Since $h$ obtained in the final part of the proof of Theorem 6.6 is a $\mathcal{T}_{c}$-retraction. Actually, we have obtained an interesting conclusion, namely the random Brouwer fixed point theorem is equivalent to the fact that $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$!
@@
By a completely similar method used in Theorem 6.6, one can easily prove that Corollary 6.5 is equivalent to the $\mathcal{T}_{\varepsilon,\lambda}$-version of random Brouwer fixed point theorem.
Acknowledgements The authors would like to thank Professor Hong-Kun Xu for his valuable comments on this work.
Funding This work was supported by the National Natural Science Foundation of China (Grant No.12371141) and the Natural Science Foundation of Hunan Province of China (Grant No.2023JJ30642).
Data availability. Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Declarations
Conflict of interest There is no conflict of interest.

### v5

@@ 6 The equivalence between the random Brouwer fixed point theorem and random Borsuk theorem
It is easy to see that both $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ and $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ are $\sigma$-stable. If $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is a $\mathcal{T}_{c}$-$r$-image of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, then there are a $\sigma$-stable $\mathcal{T}_{c}$-continuous mapping $h:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ and a $\sigma$-stable $\mathcal{T}_{c}$-continuous mapping $g:L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $h\circ g$ is the identity on $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$. Define $f:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ by f(x)=g(h(x))f(x)=g(-h(x)) for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, then $f$ is a $\sigma$-stable and $\mathcal{T}_{c}$-continuous mapping, and hence there exists $z\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ by Theorem 2.1 such that $f(z)=z$, that is to say, z=g(h(z))z=g(-h(z)), so $h(z)=(h\circ g)(-h(z))=-h(z)$. Thus $h(z)=0\in L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$, a contradiction. ∎
Since a $\mathcal{T}_{\varepsilon,\lambda}$-$r$-image and a $\mathcal{T}_{\varepsilon,\lambda}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ must be a $\mathcal{T}_{c}$-$r$-image and a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, respectively, we can immediately obtain Corollary 6.5 below, which includes [12, Lemma 3.5] as a special case.
@@
Theorem 6.4 is equivalent to the random Brouwer fixed point theorem for the random closed unite ball $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (and hence also equivalent to Theorem 2.1 by Proposition 6.1).
@@
Suppose that there exists a $\sigma$-stable and $\mathcal{T}_{c}$-continuous mapping $f:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $f$ has no fixed point. Define $e:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F})$ by $e(x)=\|f(x)-x\|$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, and denote $\eta=\bigwedge\{e(x):x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\}$. Then by [27, Lemma 3.4] there exists $x_{0}\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ such that $\eta=\|f(x_{0})-x_{0}\|$ since $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\}$ is random sequentially compact and $\sigma$-stable and since $e$ is $\sigma$-stable and random sequentially continuous. Then we have $\eta>0$, since otherwise, $\eta=0$ yields $f(x_{0})=x_{0}$, which contradicts the assumption that $f$ has no fixed point.
Now, let $A\in\mathcal{F}$ be an arbitrary chosen representative of $(\eta>0)$. Then $P(A)>0$. Since $\theta\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ and $f:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ is $\sigma$-stable, we know that $f$ has the local property, namely, $\tilde{I}_{C}f(x)=\tilde{I}_{C}f(\tilde{I}_{C}x)$ for any $C\in\mathcal{F}$ and $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$: in fact, since
@@
we have $\tilde{I}_{C}f(\tilde{I}_{C}x)=\tilde{I}_{C}f(x)$. In particular, $\tilde{I}_{A}f(\tilde{I}_{A}x)=\tilde{I}_{A}f(x)$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. Further, let
@@
and define $f_{A}:G_{A}\rightarrow G_{A}$ by $f_{A}(\tilde{I}_{A}x)=\tilde{I}_{A}f(x)$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. Then $f_{A}$ is well defined. It is clear that $G_{A}$ is a $\sigma$-stable, $\mathcal{T}_{c}$-closed, a.s. bounded and $L^{0}(\mathcal{F})$-convex subset of $L^{0}(\mathcal{F},R^{d})$ and $f_{A}$ is also $\sigma$-stable and $\mathcal{T}_{c}$-continuous.
Let $(A,A\cap\mathcal{F},P_{A})$ be the probability space with $P_{A}:A\cap\mathcal{F}\rightarrow[0,1]$ defined by $P_{A}(A\cap C)=\frac{P(A\cap C)}{P(A)}$ for any $C\in\mathcal{F}$. Further, identifying $G_{A}$ with $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$, where $\mathcal{F}_{A}=A\cap\mathcal{F}$. Then $f_{A}$ can be also identified with a mapping from $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$ to $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$. Again observing $\|f(x)-x\|>0$ on $A$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$, namely, $\|f_{A}(x)-x\|>0$ on $A$ for each $x\in G_{A}$, we can, without loss of generality, assume that $A=\Omega$. Then $\|f(x)-x\|>0$ on $\Omega$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ (since otherwise, we can consider $f_{A}$ and $L^{0}(\mathcal{F}_{A},\bar{B}_{1}(\mathbb{R}^{d}))$ to replace $f$ and $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$).
Finally, let $L_{x}=\{\lambda\in L^{0}(\mathcal{F}):\|x+\lambda(f(x)-x)\|=1\}$ for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ and $\lambda(x)=Min~L_{x}:=$ the smallest element of $L_{x}$. In fact, $L_{x}$ has only the two elements, satisfying the equation
@@
where $\langle x,y\rangle$ stands for the $L^{0}$-inner product of $x$ and $y$ in $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Similar to the root-finding formula of a quadratic equation with one variable, one can have
@@
This define a mapping $\lambda:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F})$ and it is not difficult to verify that $\lambda$ is still $\sigma$-stable and $\mathcal{T}_{c}$-continuous. By noting that $\langle x,x-f(x)\rangle=\|x\|^{2}-\langle x,f(x)\rangle$ and the random Schwartz inequality: $|\langle x,f(x)\rangle|\leq\|x\|\|f(x)\|$, we have
@@
for any $x\in L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$. In particular, when $x\in L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$, namely, when $\|x\|=1$, one has $\lambda(x)=0$. Again, define $h:L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))\rightarrow L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ by $h(x)=x+\lambda(x)(f(x)-x)$. Then $h$ is a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$ onto $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$, which contradicts Theorem 6.4. ∎
@@
The map $h$ obtained in the final part of the proof of Theorem 6.6 is a $\mathcal{T}_{c}$-retraction. Actually, we have obtained an interesting conclusion, namely, the random Brouwer fixed point theorem is equivalent to the fact that $L^{0}(\mathcal{F},\partial\bar{B}_{1}(\mathbb{R}^{d}))$ is not a $\mathcal{T}_{c}$-retraction of $L^{0}(\mathcal{F},\bar{B}_{1}(\mathbb{R}^{d}))$!
@@
By a completely similar method as used in Theorem 6.6, one can easily prove that Corollary 6.5 is equivalent to Proposition 2.3.
7 Commentaries on the state of the Schauder conjecture
Let $X$ be a Hausdorff linear topological space and $C$ be a compact convex subset of $X$. Then the famous Schauder conjecture asks if every continuous mapping from $C$ to $C$ has a fixed point. When $X$ is locally convex, the answer is affirmative, that is the famous Tychonoff fixed point theorem [54]. When $X$ is not locally convex, Cauty [8] and Ennassik and Taoudi [14] claimed that they had solved the conjecture in an affirmative way, however Mańka [39] and Yu [56] pointed out that both Cauty [8] and Ennassik and Taoudi [14] contained a gap, namely, the Schauder conjecture is still open. Although the works of this paper and [27] also involve the fixed point problems in locally noconvex spaces, for example, an $RN$ module endowed with the $(\varepsilon,\lambda)$-topology is just a metrizable locally noconvex linear topological space in general, our works are not affected by the results in [8] and [14] since our works depend on the $\sigma$-stability of the $\mathcal{T}_{\varepsilon,\lambda}$-continuous mappings and the theory of random sequentially compact $L^{0}$-convex sets rather than any works in [8] and [14]. Besides, the main challenge from [27] up to this paper lies in overcoming noncompactness since a random sequentially compact set is generally noncompact, in fact, our works also tread a new path in the study of the topological fixed point theory in $RN$ modules.
Acknowledgements The authors would like to thank Professors Hong-Kun Xu, Lixin Cheng and George Xianzhi Yuan for their valuable suggestions on this work.